Use Viete's method to get one real solution of 9x3-9x=4 and retain all roots in order to express your solution in terms of “radicals” (rather than using a decimal approximation). Simplify your solution by radicals as far as you can.
Answer #1

et L, B, Y be roots cao 0 9x3 - 91=4 9x39x-4=0 using vietes method & + ßty=0 XB+BY+Y2=-9/9 LBY Alg x² - x -4 = 0 x² + px + q